# Right Place, Right Time! Dynamizing Topological Graphs for Embodied Navigation

### v3

@@ 9.1 LGX Inference
User Prompt - “Which object from $<$OBJECT_LIST$>$ should I go towards to find a new portable object? Reply in ONE word.”
The $<$OBJECT_LIST$>$ here contains a set of objects that have been detected using YOLO-v8, as proposed in LGX. Additionally, we if a portable object is present at a certain node at a given timestep, we add it to this list.
The LLM then predicts a target object from the list, which is mapped to an adjacent node using our customized MP3D functions.
For the memory-enhanced LGX case, we modify the System Prompt, with memory. We additionally ask -
@@
1. $<$OBJECT_LIST$>$: ACTION
2. $<$OBJECT_LIST$>$: ACTION
@@
“Which object from $<$OBJECT_LIST$>$ should I go towards to find a new portable object? Reply in ONE word.”
10 Code and Dataset

### v4

@@ 8.3 TAS-LLM
User Prompt - “Which object from <<OBJECT_LIST>> should I go towards to find a new portable object? Reply in ONE word.”
The <<OBJECT_LIST>> here contains a set of objects that have been detected using YOLO-v8, as proposed in LGX. Additionally, we if a portable object is present at a certain node at a given timestep, we add it to this list.
The LLM then predicts a target object from the list, which is mapped to an adjacent node using our customized MP3D functions. The agent then moves to that node.
For the TAS-LGX case, we modify the System Prompt, with memory as presented in Section 3.3. We additionally ask -
@@
1. <<OBJECT_LIST>>: ACTION
2. <<OBJECT_LIST>>: ACTION
@@
“Which object from <<OBJECT_LIST>> should I go towards to find a new portable object? Reply in ONE word.”
Note that we maintain a horizon to avoid token overflow, where the earliest appended observation is removed when the context length limit is reached.
9 Video, Code and Dataset